The U.S. Navy has proposed a system, called ELF (extra long frequency) for communicating with submerged submarines. The system uses radio waves with a frequency of 76 s-1. What is the wavelength of this radiation in meters?

The yellow light emitted by the sodium lamps in streetlights have wavelengths of 589.6 nm and 589.0 nm. What is the frequency of the 589.0 nm light? What is the energy of 1 mole of photons of yellow light with a wavelength of 589.0 nm?
1
Expert's answer
2015-11-03T11:56:22-0500
λ = с/ν = 299792458 m s-1/76 s-1 = 3944637 m
ν = с/λ = 299792458 m s-1/589.0*10-9 m = 509.0 MHz
E = hcNA/λ = 1.98644568*10-25*6.022*1023 mol-1/589.0*10-9 m = 203.1 kJ mol-1